### v1

# arXiv:2603.02685v1[cond-mat.supr-con] 03 Mar 2026

Abstract
As an unconventional superconducting system capable of reaching 40K under ambient pressure, the La3Ni2O7 film superconductor has become a recent focal point in the field of superconductivity, demanding further theoretical exploration of possible pairing mechanisms. In this work, we employ the fluctuation exchange (FLEX) approximation to systematically analyze the superconducting properties of the two-orbital, two-site model for the La3Ni2O7 film in the weakly correlated regime, focusing on its dependence on hole-doped concentration. Through a more detailed examination of the Fermi surface topology, we find that when a $\delta$ pocket composed of the $d_{z^{2}}$ orbital emerges near the $\Gamma$-point, its nesting with the $\gamma$-pocket, along with the nesting between the $\alpha$- and $\beta$-pockets, leads to a mutual enhancement of $s_{\pm}$-wave pairing at the corresponding wave vector. Furthermore, we propose that this nesting-driven enhancement of spin-fluctuation-induced pairing may be a viable mechanism to enhance superconductivity.
I Introduction.
The mechanism of high-temperature superconductivity has long been a central research topic in condensed matter physics. In recent years, following the discovery of cuprate and iron-based superconductors, the emergence of the nickel-based superconductor family LDF2019 ; Wang327_Nature2023 has attracted wide attention. Among them, Ruddlesden-Popper (RP) phase nickel-based superconductors, exemplified by the La3Ni2O7 compound, can achieve a superconducting transition temperature of up to 80K under high pressure conditions Wang327_Nature2023 . During the past two years, significant progress has been made in both experimental characterization hou2023emergence ; Zhang_2024_HighTc ; Xie_2024_MagneticExchange ; Li_2024_ElectronicCorrelation ; Wang_2024_La2PrNi2O7_Nature ; dan_2024 ; wang2024pressure ; cai_2024 ; liu_2024 and theoretical studies YaoPRL2023 ; Yang_2023_sPlusMinus ; Liu_2023_sPlusMinus ; Yang_2023_ValenceBonds ; Qin_2023_HighTc ; zhang_2023 ; pupha_l2023 ; Christiansson_2023 ; pupha_l2023 ; liao2023electron ; liao_2024 ; zhang2024electronic ; zhang2023trends ; Tian_2024_Correlation ; Oh_2024_StrongPairing ; Fan_2024_BilayerCoupling ; RyeePRL2024CMDFT ; wang2024electronic ; chen2024electronic ; lu2024interlayer ; sakakibara2024 ; jiang2024high ; zhang2024strong ; Zheng_2025_TwoOrbital ; jiang2025theory ; duan2025orbital ; xi2025transition ; shi2025theoretical ; ouyang2025two ; maier2025interlayer ; wang2025electronic ; puphal2025superconductivity of high-pressure nickel-based superconductors. One of the central questions in the theoretical study of nickel-based superconductors is how to further enhance their superconductivity, enabling them to operate at lower pressures and higher temperatures. The shape of the Fermi surface plays a critical role in determining superconducting behavior Wang327_Nature2023 ; Yang_2023_sPlusMinus ; Liu_2023_sPlusMinus ; Yang_2023_ValenceBonds . Importantly, the Fermi surface is sensitive to variations in lattice constants, electron filling, and interaction strength wu2024ac ; jiang2024high ; huo2025modulation ; wang2025electronic . Theoretically, many research groups have explored various strategies such as chemical doping, applying pressures to tune the Fermi surface geometry and the density of states at the Fermi level, trying to enhance superconductivity wang2023observation ; wu2024ac ; rhodes2024structural ; geisler2024structural ; chen2025charge ; huo2025modulation .
Recently, the research team led by Z.Y.Chen achieved a breakthrough by successfully growing a La2.85Pr0.15Ni2O7 thin-film superconductor on an SrLaAlO4 substrate under ambient pressure, with a superconducting transition temperature reaching about 40K Ko_LNO327AmbientSC ; Zhou_2024_AmbientSC , stimulating extensive experimental and theoretical discussions regarding their superconducting mechanism liu2025superconductivity ; XJZhou_ARPES_327 ; li2025photoemission . The theoretical study by the authors yue2025correlated reveal this complex heterostructure is a moderately correlated system far from doped Mott insulator regime. Our results suggested that the superconductivity in ambient pressure La3Ni2O7 films may originate from spin-fluctuation-induced $s_{\pm}$-wave pairing.
However, the significance of the ambient-pressure La3Ni2O7 thin film superconductor extends far beyond this single material. Its study may not only guide the discovery of nickel-based superconductors with higher superconducting transition temperatures under ambient conditions, but also suggests that the structure of La3Ni2O7 thin films can be tuned by selecting different substrates osada2025strain . This, to some extent, alleviates the experimental difficulty of doping in nickel-based systems and provides a new platform for gaining deeper insights into the superconducting pairing mechanism hao2025superconductivity ; wang2025recent .
We find that, compared with the bulk La3Ni2O7 compound, the La2.85Pr0.15Ni2O7 thin-film structure exhibits a larger inter-layer Ni–Ni distance within each bilayer, which directly reduces the inter-layer hopping $t_{\perp}^{z}$ between the $d_{z^{2}}$ orbitals yue2025correlated . A smaller $t_{\perp}^{z}$ brings the bonding and anti-bonding states of the $d_{z^{2}}$ orbital closer in energy and induces a tendency toward the emergence of a new Fermi surface. If the $c$-axis lattice constant can be further expanded experimentally, it would likely be favorable for enhancing the superconductivity of nickel-based systems.
In this work, we employ the fluctuation exchange (FLEX) approximation to investigate the La3Ni2O7 thin film system in the weak coupling regime. By analyzing the Fermi surface nesting mechanism, we predicted a possible scenario that could enhance superconductivity in this system. Specifically, we first discuss the key features of the FLEX-corrected Fermi surface and spin correlation functions. We then compute and analyze the leading eigenvalue $\lambda$ of the linearized Eliashberg equation and its corresponding gap function on the Fermi surface, exploring possible enhance mechanisms of pairing based on the numerical behavior of $\lambda$. Finally, we summarize and discuss our findings.
II Method
Figure 1: (a), The schematic crystal structure of half-UC thin film of La3Ni2O7 using the lattice parameters of La2.85Pr0.15Ni2O7 grown on SrLaALO4 substrate Zhou_2024_AmbientSC , where nickel atoms form octahedral coordination with surrounding oxygen atoms. For a better illustration, we put a exaggeratedly small vacuum than the one used in calculation above the thin film. (b), The band structure of the half-UC system. The orange lines show the DFT bands, green lines the TB bands. We also show the projected bands of the $e_{g}$ orbitals, with the red (blue) dots the $d_{x^{2}-y^{2}}$ ($d_{z^{2}}$) orbital.
In this work, we simplify the complicated La2.85Pr0.15Ni2O7/SrLaAlO4 heterostructure to an half-unit-cell (half-UC) La3Ni2O7 with the experimental parameters, including the lattice constants, atomic spacing (see Ref. yue2025correlated ). The minimal two-site two orbital ($d_{x^{2}-y^{2}}$ and $d_{z^{2}}$ orbitals per Ni sites) model Hamiltonian is also adopted. We use $A,B$ to label the up layer and down layer and $x,z$ to label the orbital $d_{x^{2}-y^{2}}$ and $d_{z^{2}}$. So, the orbital basis in site representation can be expressed as $\Psi=[Ax,Az,Bx,Bz]^{T}$. The same Hamiltonian as in previous studies is employed yue2025correlated , which can be written in momentum space, reads
$H=H_{tb}+H_{int},$ (1)
where the tight-binding (TB) model part reads
$H_{tb}(k)=\begin{pmatrix}T_{k}^{x}&V_{k}&T^{{}^{\prime}x}_{k}&V_{k}^{{}^{\prime}}\\ V_{k}&T_{k}^{z}&V_{k}^{{}^{\prime}}&T^{{}^{\prime}z}_{k}\\ T^{{}^{\prime}x}_{k}&V_{k}^{{}^{\prime}}&T_{k}^{x}&V_{k}\\ V_{k}^{{}^{\prime}}&T^{{}^{\prime}z}_{k}&V_{k}&T_{k}^{z}\end{pmatrix}$. (2)
Here,
$T_{k}^{x/z}$ $=2t_{1}^{x/z}(\cos k_{x}+\cos k_{y})+2t_{4}^{x/z}(\cos 2k_{x}+\cos 2k_{y})$
$+2t_{5}^{x/z}(\cos 3k_{x}+\cos 3k_{y})+4t_{2}^{x/z}\cos k_{x}\cos k_{y}$
$+\epsilon_{x/z},$ (3)
$T^{{}^{\prime}x/z}_{k}=t_{\perp}^{x/z}+2t_{3}^{x/z}(\cos k_{x}+\cos k_{y}),$ (4)
$V_{k}=2t_{3}^{xz}(\cos k_{x}-\cos k_{y})+2t_{5}^{xz}(\cos 2k_{x}-\cos 2k_{y}),$ (5)
and
$V_{k}^{{}^{\prime}}=2t_{4}^{xz}(\cos k_{x}-\cos k_{y}),$ (6)
The hopping parameters can be found in Table 2 of our previous work yue2025correlated . The TB band structure is shown by the green lines in Fig. 1(b), which agrees well with the DFT one (orange line). The interaction Hamiltonian takes the Hubbard-Kanamori form, which reads
$H_{\text{int}}=$ $U\sum_{i\text{, }\mu}n_{i\mu\uparrow}n_{i\mu\downarrow}+U^{\prime}\sum_{i\text{, }\mu<\nu\text{, }\sigma}n_{i\mu\sigma}n_{i\nu\bar{\sigma}}$
$+(U^{\prime}-J)\sum_{i\text{, }\mu<\nu\text{, }\sigma}n_{i\mu\sigma}n_{i\nu\sigma}$
$-J\sum_{i\text{, }\mu\neq\nu}c^{\dagger}_{i\mu\uparrow}c_{i\mu\downarrow}c^{\dagger}_{i\nu\downarrow}c_{i\nu\uparrow}$
$+J\sum_{i\text{, }\mu\neq\nu}c^{\dagger}_{i\mu\uparrow}c^{\dagger}_{i\mu\downarrow}c_{i\nu\downarrow}c_{i\nu\uparrow}$ (7)
Here, $U$ ($U^{\prime}$) is the onsite intra-orbital (inter-orbital) Coulomb interaction, and $J$ the onsite Hund coupling and pair hopping. We assume a conventional choice for $U^{\prime}$ with $U^{\prime}=U-2J$. Furthermore, $J=U/6$ is chosen throughout this work.
Figure 2: The Fermi surface of $n=1.3$ (a) and $n=1.42$ (b). The spin susceptibility of $n=1.3$ (c) and $n=1.42$ (d). The four Fermi pocket named as $\delta,\alpha,\beta,\gamma$ are labeled in (b). The nesting wave vectors are marked in the corresponding figures.
FLEX is a conserving quantum many-body approximation method that self-consistently incorporates spin and charge fluctuations by summing infinite-order bubble and ladder diagrams in the self-energy calculation bickers1989conserving ; bickers1991conserving ; yanase2003theory ; kuroki2008unconventional ; witt2021efficient ; bjornson2021orbital . While FLEX is a perturbative approach, it has been proven highly effective in capturing the pairing symmetry driven by Fermi surface nesting. In this work, we employ the FLEX method to calculate the band renormalization of half-UC film La3Ni2O7 system induced by interactions within the weak coupling regime. Here, only the particle-hole channel is considered. Its effective interaction is given by
$V(q)=$ $\frac{3}{2}U^{S}\chi^{S}(q)U^{S}+\frac{1}{2}U^{C}\chi^{C}(q)U^{C}-\frac{3}{4}U^{S}\chi^{0}(q)U^{S}$
- $\frac{1}{4}U^{C}\chi^{0}(q)U^{C}+\frac{3}{2}U^{S}-\frac{1}{2}U^{C}$ (8)
We can write the matrix element of $U^{S},U^{C}$ in site representation as
$U^{S}_{ijkl}=\begin{cases}U\text{\ ,}&i=j=k=l\\ U^{\prime}\text{, }&i=k\neq j=l\\ J\text{\ ,}&i=j\neq k=l\\ J^{\prime}\text{, }&i=l\neq j=k\end{cases},$ (9)
$U^{C}_{ijkl}=\begin{cases}U\text{\ ,}&i=j=k=l\\ -U^{\prime}+J\text{, }&i=k\neq j=l\\ 2U^{\prime}-J\text{\ ,}&i=j\neq k=l\\ J^{\prime}\text{, }&i=l\neq j=k\end{cases}$ (10)
Please notice that only intra-layer elements are non-zero. Here, $\chi^{0}(q)$ is the bare susceptibility that reads $\chi^{0}_{ijkl}(q)=-\frac{T}{N}\sum_{k}G_{ki}(k)G_{jl}(k)$. $G_{ij}(k)$ is the Green’s function calculated by Dyson’s equation. $\chi^{S}(q),\chi^{C}(q)$ are the spin and charge susceptibilities, which can be calculated by
$\chi^{C/S}(q)=\frac{\chi^{0}(q)}{1\pm U^{C/S}\chi^{0}(q)}.$ (11)
In this work, we use a modified code of FLEX by Witt witt2021efficient . In our calculations, we adopt a k-grid as large as $64\times 64$ , and the temperature is fixed at $T=0.001eV$. To ensure the validity of the FLEX approach, we restrict our discussion to cases where $U$ is relatively small. After achieving self-consistent convergence in FLEX, the obtained physical quantities are used to solve the linearized Eliashberg equation to extract information about superconducting pairing. The self-consistent gap equation to be solved is written as
$\lambda\Delta_{il}(k)=\frac{T}{N}\sum_{q}\sum_{jk}V_{ijkl}(q)F_{jk}(k-q)$ (12)
where $F(k)=-G(k)\Delta(k)G^{T}(\mathbf{-k})$ is the anomalous Green’s function, The closer the eigenvalue $\lambda$ of the self-consistent equation approaches $1$, the larger the instability towards the superconducting state with a certain pairing symmetry. In previous works, we found that pairing in this system is always predominantly spin-singlet yue2025correlated . Therefore, we focus primarily on spin-singlet pairing, namely
$V(q)=\frac{3}{2}U^{S}\chi^{S}(q)U^{S}-\frac{1}{2}U^{C}\chi^{C}(q)U^{C}+\frac{1}{4}(3U^{S}+U^{C})$ (13)
The largest eigenvalue of the linearized Eliashberg equation is efficiently computed using the power method. The pairing symmetry is analized from the gap function $\Delta(\mathbf{k,i\omega_{0}})$ of this system, where $\omega_{0}$$\pi T$ is the lowest Matsubara fermi frequencies.
Figure 3: (a) The $\lambda$ as a function of electron filling $n$ at $U=1.5$, with the red line corresponding to $d_{xy}$-wave symmetry and the black line the $s$-wave symmetry. (b)-(c) The gap function of the linearized Eliashberg equation within $s$- and $d_{xy}$-wave symmetry, respectively, at $n=1.3$. (d) The $\lambda$ in $s$-wave symmetry as a function of $U$, with the red line corresponding to $n=1.42$ and the black line $n=1.3$. (e)-(f) The gap function of the linearized Eliashberg equation within $s$- and $d_{xy}$-wave symmetry , respectively, at $n=1.42$.
III Results
Since the previous computational results for the La3Ni2O7 superconducting film indicate that the electron filling $n$ should lie within the hole doped regime yue2025correlated , we first choose two different hole doping levels for discussion. To ensure the reliability of FLEX, we set $U=1.5$ throughout. Fig. 2 (a) and (b) illustrate the self-consistent Fermi surfaces (FS) for $n=1.3$ and $n=1.42$. The FS is composed of three electron pockets and one hole pocket, labeled $\delta,\alpha,\beta,$ and $\gamma$ in order of increasing distance to the $\Gamma$ point. As hole doping increases, the hole-like $\gamma$ pocket shrinks; at $n=1.3$, it is on the verge of vanishing. Fig. 2(c) and (d) show the largest eigenvalues of the spin susceptibility corresponding to $n=1.3$ and $n=1.42$, respectively. At $n=1.3$, the spin susceptibility exhibits four distinct series of peaks, which are marked in Fig. 2(a). Each of these peaks corresponds to a specific nesting condition in the Fermi surface: (1) $Q_{1}$ has two possible nesting contributions, one from the $\delta$ pocket to the $\gamma$ pocket and the other from the $\alpha$ pocket to the $\beta$ pocket. (2) $Q_{2}$ corresponds to the nesting from the $\alpha$ pocket to the $\gamma$ pocket. (3) $Q_{3}$ corresponds to the nesting within the $\gamma$ pocket to itself. (4) $Q_{4}$ corresponds to the nesting from the $\beta$ pocket to the $\gamma$ pocket. In contrast to the multi-peak structure at $n=1.3$, the spin susceptibility at $n=1.42$ is dominated by peaks at $Q_{1}$ and $Q_{4}$, with the $Q_{1}$ mode being particularly pronounced. This enhancement arises from two factors. First, the $\gamma$ pocket expands at $n=1.42$, and its curvature along the diagonal direction flattens, leading to a more extended tangential region with the $\beta$ pocket. Second, the nesting between the $\alpha$ and $\delta$ pockets improves due to better geometric congruence (i.e., the nested segments are nearly parallel). These factors cooperatively strengthen the spin fluctuations at $Q_{1}$. Concurrently, the shrinkage and increased curvature of the $\gamma$ pocket suppress the scattering channels at $Q_{2}$ and $Q_{3}$.
Figure 4: (a)(b) The FS which the $\delta$ pocket is removed at $n=1.3$ and $n=1.42$. (d)(e) The gap function of $s$-wave symmetry calculated by FS like (a) and (b). (c)(f) The $\lambda$ as a function of U($n$), n = 1.42 in (c) and U = 1.5 in (f).
Fig. 3(a) shows the variation of the eigenvalues $\lambda$ for s-wave and d-wave solutions of the linearized Eliashberg equation as a function of electron filling $n$. It can be seen that in the over-doped hole-doped region ($n<1.3$), both s-wave and d-wave $\lambda$ values remain relatively small and exhibit competition, with the instability d-wave pairing having a slight advantage. However, in the region of lower hole doping, the s-wave eigenvalue $\lambda_{s}$ increases significantly with increasing $n$ and then forming a dome-like structure near $n=1.42$. This phenomenon can be understood by considering the FS and spin susceptibility as shown in Fig. 2, along with the gap function. Fig. 3(d) shows the variation of $\lambda_{s}$ as a function of interaction strength $U$ for two different doping levels. As $U$ increases, $\lambda_{s}$ exhibits a steady growth. Fig. 3(b) and (c) show the gap functions on the Fermi surface for s-wave and d-wave symmetries at $n=1.3$, obtained after the power method converges and is already transformed into the band representation. Among the four types of nesting at this electron filling, only $Q_{3}$ (see Fig. 2) occurs within the intra-band, providing spin fluctuations that favor $d_{xy}$-wave pairing. In contrast, the other three inter-band nesting processes contribute to the possibility of $s_{\pm}$-wave pairing. From the gap function analysis, although the $\delta$ pocket has almost vanished, the relatively flat band near the $\Gamma$-point still generates a noticeable signal. This suggests that in the over-doped hole-doped region, there exists a competition between $s_{\pm}$-wave pairing, primarily driven by $Q_{1}$ and $Q_{4}$, and $d_{xy}$-wave pairing, dominated by $Q_{3}$. Fig. 3(e) and (f) show the gap functions for the s-wave and d-wave symmetries at $n=1.42$, where this doping level is located at the dome of the $\lambda_{s}(n)$ which is closest to 1. Compared to the case of $n=1.3$, the Fermi surface at this electron filling exhibits a larger density of states on the $\delta$ pocket, which matches better in size with the $\gamma$ pocket. This suggests that the significant enhancement of superconductivity at this doping level is driven by the strengthened spin fluctuations at the $Q_{1}$ wave vector, which favor pairing. However, for lower hole doping, although the shape matching between the $\delta$ pocket and the $\gamma$ pocket may improve the nesting, the density of states in the $\delta$ pocket decreases due to departure from the flat band near the $\Gamma$ point. Consequently, the spin fluctuations at $Q_{1}$ do not continue to increase, which may explain the presence of the dome structure at $\lambda_{s}(n)$.
In the previous discussion, we observed that in the weak coupling regime, the s-wave superconductivity in this system may be contributed by four types of nesting and the $\delta$ pocket appears to have a significant impact on superconductivity through the nesting with the $Q_{1}$ wave vector. To further investigate the impact of the $\delta$ pocket on superconductivity, we performed a computational experiment by artificially suppressing the Green’s function spectral weight around the $\Gamma$ point in the Eliashberg calculation. This allows us to isolate the specific contribution of the $\delta$ pocket to the superconducting instability. After artificially removing the $\delta$ pocket, the Fermi surfaces for $n=1.3$ and $n=1.42$ are shown in Fig. 4(a) and (b). We compared $\lambda_{s}$ at $n=1.42$ in cases where the $\delta$ pocket is present and absent, with the results shown in Fig. 4(c). The gap functions after removing the $\delta$ pocket are displayed in Fig. 4(d) and (e). The original $\lambda_{s}$ is consistently higher than that of the case where the $\delta$ pocket is removed for all values of $U$, indicating that although the nesting between the $\alpha$ pocket and $\beta$ pocket, as well as that between the $\gamma$ pocket and $\beta$ pocket (i.e. spin fluctuation of wave vector $Q_{1}$ and $Q_{3}$), still induce the $s_{\pm}$-wave pairing. However, in the presence of the $\delta$ pocket, the more stable $s_{\pm}$-wave pairing occurs between the $\delta$ pocket and the $\gamma$ pocket. Fig. 4(f) shows the evolution of $\lambda_{s}$ with $n$ in both cases, demonstrating that this enhancement remains effective across a broad hole-doped regime. Moreover, after removing the $\delta$ pocket, the dome structure previously observed at $n=1.42$ disappears, confirming that the optimal $T_{c}$ is dictated by the evolution of the $d_{z^{2}}$-derived Fermi surface near the $\Gamma$ point.
IV Discussion and Conclusions.
In this work, we used the FLEX method to study a previously proposed model Hamiltonian for the ambient pressure La3Ni2O7 film superconductor, which is a two-site two-orbital Hubbard model. We investigated the superconducting properties of this model in the weakly correlated, hole-doped region. We conclude that the complex spin fluctuations induced by repulsive Coulomb interactions can facilitate $s_{\pm}$-wave pairing on the $d_{z^{2}}$ orbitals, which aligns with the recent theoretical findings by Gao gao2025theoretical . By examining the features of the Fermi surface and correlation functions, we discussed the impact of four spin susceptibility peaks caused by five different nesting mechanisms on superconductivity. We found that a non-trivial $\delta$ pocket in the studied band structure significantly enhances the tendency toward $s_{\pm}$-wave superconductivity. Furthermore, we analyzed the consistency between the Fermi surface, spin susceptibility, superconducting instability indicator $\lambda$, and the gap function as they evolve with electron filling, $U$. Finally, we confirmed this superconductivity enhancing mechanism by artificially tuning the presence or absence of the $\delta$ pocket.
The $\delta$ band in realistic thin-film of La3Ni2O7 ever reported tends to be away from the Fermi level. However, by slightly modifying the lattice structure through in-plane strain LHY2507 , doping or substrate substitution, the $\delta$ pocket could be ”brought back” to the Fermi surface in new material realizations. If achieved in the future experiment, the presence of the $\delta$ pocket and its enhancement mechanism proposed in this study could lead to a higher superconducting transition temperature in nickel-based superconductors.
Acknowledgements. — Changming Yue acknowledges support from the National Natural Science Foundation of China (12474231), Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401004). Jian-Jian Miao is supported by NSFC (No. 12404171), the Guangdong Project (Grant No. 2024QN11X176).